What concentration results when 1.7 L of 0.45 mol/L Ba (OH)2 is mixed with 600 mL of 1.70 mol/L Ba(OH)2?


1
Expert's answer
2022-05-31T15:14:04-0400

Solution:

Liters of solution 1 = 1.7 L

Moles 1 of Ba(OH)2 = (0.45 mol/L) × (1.7 L) = 0.765 mol

Liters of solution 2 = (600 mL) × (1 L / 1000 mL) = 0.6 L

Moles 2 of Ba(OH)2 = (1.70 mol/L) × (0.6 L) = 1.020 mol


Total moles of Ba(OH)2 = Moles 1 + Moles 2 = 0.765 mol + 1.020 mol = 1.785 mol

Total liters of solution = Liters of solution 1 + Liters of solution 2 = 1.7 L + 0.6 L = 2.3 L


Molarity = Moles of solute / Liters of solution

Therefore,

Molarity of Ba(OH)2 solution = (1.785 mol) / (2.3 L) = 0.776 mol/L

Molarity of Ba(OH)2 solution = 0.776 mol/L


Answer: 0.776 mol/L